OpenStudy (anonymous):

Which equations show a direct variation between x and y when solved for y? Choose all answers that are correct. A. y-9=2x+9 B. y= 4/7x C.4+y=1/3x + 4 D.y=3-x

OpenStudy (anonymous):

@mathmath333

OpenStudy (alexandervonhumboldt2):

direct variation has form y=kx

OpenStudy (anonymous):

What does the k stand for?

OpenStudy (mathmath333):

k can be any number other than 0

OpenStudy (alexandervonhumboldt2):

k is just a coefficient

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

dont forget to simplify all equations: there are more than 1 answer.

OpenStudy (anonymous):

So would D.y=3-x be one of them because x=0?

OpenStudy (alexandervonhumboldt2):

nooooo

OpenStudy (alexandervonhumboldt2):

D has form y=kx+b and we need y=kx

OpenStudy (anonymous):

Well B and D are kinda like the same thing so can the be crossed out

OpenStudy (alexandervonhumboldt2):

nooooB is actually corect

OpenStudy (alexandervonhumboldt2):

y= 4/7x k=4/7. so it satisfies y=kx form

OpenStudy (anonymous):

Ohhh

OpenStudy (alexandervonhumboldt2):

try finding a second answer now there is one more

OpenStudy (anonymous):

Would it be C? Because 1/3x or 1/3

OpenStudy (alexandervonhumboldt2):

yeah C is correct as well.

OpenStudy (alexandervonhumboldt2):

@pooja195 if you are global. please do not report me. I did NOT gave a direct answer. have a nice day.

OpenStudy (anonymous):

Its fixed thanks for the help @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

:) yw